OpenStudy (anonymous):

(Complex Numbers) Find each quotient: (2-3i)/(2+3i)

OpenStudy (anonymous):

do you know what a conjugate is?

OpenStudy (anonymous):

yes, I multiplied the numerator and denominator by: (2-3i) but I don't see how my answer is going to be -5/13-12/13i. So I don't know what I am doing wrong.

OpenStudy (anonymous):

That's what the answer key says it's suppose to be.

OpenStudy (anonymous):

that answer looks good to me

OpenStudy (anonymous):

How do I get there? I'm at (4-12i+9i^2)/(4-9i^2)

OpenStudy (anonymous):

the denominator is \(a^2+b^2\) which in your case is \(2^2+3^2=4+9=13\)

OpenStudy (anonymous):

I^2=-1

OpenStudy (anonymous):

oh i see, you have the right answer, just forget that \(i^2=-1\)

OpenStudy (anonymous):

I got: (3+3i) after I did the rest, and that's still not in the fractions that the back of the book has, so how does it get to be: -5/13-12/13i ?

OpenStudy (anonymous):

I mean (-3+3i) was what I got.

OpenStudy (anonymous):

you wrote \[\frac{4-12i+9i^2}{4-9i^2}\]right?

OpenStudy (anonymous):

Yeah, and then put (-1) in place of the I^2 and got (-3+3i)

OpenStudy (anonymous):

\[\frac{4-12i+9i^2}{4-9i^2}=\frac{4-12i+9\times (-1)}{4-9\times (-1)}\]

OpenStudy (anonymous):

you get \[\frac{4-9-12i}{4+9}=\frac{-5-12i}{13}\]

OpenStudy (anonymous):

then in standard form this is \[-\frac{5}{13}-\frac{12}{13}i\]

OpenStudy (anonymous):

clear or no? i am not really sure how you got the \(3-3i\) from this

OpenStudy (anonymous):

Yeah I see what I was doing wrong right there now, I was subtracting the -1 instead of multiplying it, thank you for your help!